Session: May 2002 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Fri May 31 00:56:10 CEST 2002) Success
No problems.

Urumqi (Liu Xiang at Fri May 31 04:42:15 CEST 2002) Minor failures
Tsys is high, because of cooling problem. FS9.5.11 was used, it gave error messages: Atenna time is wrong by more than 3 seconds, antenna is not tracking. In Fact, we found the antenna was tracking for every source, so no problem.

\(Other\) (Vargiu at Fri May 31 05:41:24 CEST 2002) Minor failures
No problem during the session.

Jodrell_Bank (Alastair Gunn at Fri May 31 06:42:52 CEST 2002) Minor failures
Some VCs may have intermittently lost lock due to power supply problems. No other known problems.

Torun (Kaz Borkowski at Fri May 31 07:24:28 CEST 2002) Minor failures
BBC7 unlock-ed all the time. No other known problems.

Effelsberg (A. Kraus at Fri May 31 12:30:22 CEST 2002) Success

Noto (P. Leto at Mon Jun 3 09:27:04 CEST 2002) Success
No known problems

Westerbork (Tony Foley at Mon Jun 3 09:54:13 CEST 2002) Minor failures
Scans before 13:55 out of hour angle range and replaced with local cal. Array Phasing reasonable.

Onsala (John Conway at Mon Jun 3 14:28:24 CEST 2002) Minor failures
No known problems